if sarah has grades of 77 and 91 on her first two test. If she wants atleast 80 after her third test, what score must she make on that test?
Answer by mananth(16949)

if sarah has grades of 77 and 91 on her first two test. If she wants atleast 80 after her third test, what score must she make on that test?
..
average required out of 3 tests = 80 .
total in these three tests = 80*3 = 240
..
In 2 tests she has obtained 77+91 = 168
..
Balnce is what she has to score;
240-168=72
...
m.ananth@hotmail.ca